OpenStudy (anonymous):

How many centimeter (squared) = 1 meter (squared)?

OpenStudy (anonymous):

1cm = 1* 10^-2 m (1 cm)^2 = (1 * 10^-2m)^2 1 cm^2 = 1 * 10^-4 m^2 dividing by 10^-4 (or multiplying by 10000 :-P) 1 m^2 = 10000 cm^2

OpenStudy (jamesj):

Imagine a square with sides one meter long. Now divide each side up into 100 cm. Then you have a grid in the square 100 squares times

OpenStudy (jamesj):

100 squares. Or 100^2 = 10,000 squares. hence 1 sq meter = 10,000 sq cm

OpenStudy (anonymous):

Thank you, youre a lifesaver!